THE LINEAR SAMPLING METHOD FOR ANISOTROPIC MEDIA, PART II

FIORALBA CAKONI HOUSSEM HADDAR

Abstract. We reconsider the linear sampling method for solving the inverse scattering problem of determining the support of an anisotropic inhomogeneous medium from a knowledge of the incident and scattered time harmonic acoustic wave at fixed frequency. We extend the results of a previous paper concerning with the same problem to the case where the norm of the matrix that describes the physical properties of the medium is less than one.